# Graph Classification Gaussian Processes via Hodgelet Spectral Features

### v1

@@ 4 Conclusion and Future Directions
We have presented a GP-based classification method that can classify graphs based on vertex and/or edge features. The vertex and edge features were processed by considering their projections onto the eigenbases of the graph Laplacian and Helmholtzian, respectively. The resulting wavelet coefficients were then used to construct wavelet spectral features, which were used as input points to a standard GP classification algorithm. Further, by taking the Hodge decomposition, we have shown that it can provide improvements over the method by Opolka et al. [5], even on graph data with only vertex features, owing to the added flexibility of the model. Overall, we have demonstrated the efficacy of our approach for graph classification tasks where edge features or topological information are essential. In the future, we intend to consider extensions to higher-order networks, such as simplicial complexes, cellular complexes, and hypergraphs, which we illustrate briefly in Appendix C.
@@ Appendix A Gaussian Processes
Gaussian processes are stochastic processes that can be considered as Gaussian distributions extended to infinite dimensions. More precisely, a Gaussian process over $\mathbb{R}^{d}$ is a random function $f:\Omega\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, such that for any finite sequence of input points $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{M}\in\mathbb{R}^{d}$, the random vector $[f(\boldsymbol{x}_{1}),\ldots,f(\boldsymbol{x}_{M})]\in\mathbb{R}^{M}$ is jointly Gaussian. Gaussian processes are defined by a mean function $\mu:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and a kernel function $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, which satisfy $\mu(\boldsymbol{x})=\mathbb{E}[f(\boldsymbol{x})]$ and $\kappa(\boldsymbol{x},\boldsymbol{x}^{\prime})=\mathrm{Cov}[f(\boldsymbol{x}),f(\boldsymbol{x}^{\prime})]$, respectively, for any $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}$. We denote this relation by writing that $f\sim\mathrm{GP}(\mu,\kappa)$. Gaussian processes are employed in the Bayesian inference paradigm: (1) a Gaussian process is defined as a prior on an unknown function, and (2) combined with a likelihood and training dataset, this produces a posterior process (not necessarily a Gaussian process). In the case of classification, the likelihood is non-Gaussian. For example, the Bernoulli likelihood is convenient for binary classification and, more generally, the categorical likelihood for multi-class classification. The ensuing posterior process is intractable and needs to be approximated, which can be done using variational methods.
@@
A.1 Kernels
Gaussian process is uniquely characterised by a mean function and kernel function, which has the advantage of being simple and interpretable. The mean function is often taken to be zero (along a data normalisation), i.e. $\mu(\boldsymbol{x})=0$, for every $\boldsymbol{x}\in\mathbb{R}^{d}$, as it simplifies derivations and proofs. The kernel function is thus considered the most defining component in describing the behaviour of a Gaussian process and measures the similarities between two input points. It is important to note that a kernel function must be a symmetric and positive semi-definite function. Some popular choices are the square exponential kernel function (also called radial basis function kernel)
@@ A.1 Kernels
A.2 Computational cost
@@
Computing the posterior requires the inversion of a $N\times N$ Gram matrix, where $N$ is the number of training input points. This results in a cubic computational complexity $\mathcal{O}(N^{3})$ and quadratic memory requirements $\mathcal{O}(N^{2})$. Fortunately, sparse Gaussian process [24] have been devised to alleviate this computational burden by constructing a smaller training dataset of $M$ input points, called inducing points, where $M\ll N$.
Appendix B Wavelet Transforms
Spectral analysis is a set of tools for analysing and manipulating signals according to the eigenvalues, sometimes referred to as frequencies, of the Laplacian operator. The Fourier transform is the projection of signals onto the eigenvectors, called Fourier basis, of the Laplacian operator, thus producing Fourier coefficients. Although this provides a complete localisation in terms of frequency, it does not offer spatial localisation, and the resolution is fixed, i.e. it cannot capture localised changes in frequency. The wavelet transform provides both a balanced analysis between the spatial and spectral domains, and multi-scale resolution. The wavelet transform modulates the Fourier coefficients via functional calculus on the Laplacian operator. For graphs, this is done by applying a combination of two functions, a wavelet function and a scaling function, to the eigenvalues of the graph Laplacian. A wavelet function is a scaled and shifted variant of a mother wavelet, acting as a band-pass filter, designed to cover medium and high frequencies. A scaling function acts as a low-pass filter, capturing low frequencies.
@@ Appendix B Wavelet Transforms
The reader is referred to Hammond et al. [8] for a detailed description of the graph wavelet transform.
B.1 Mother wavelet
@@ Appendix C Extension to Higher-Order Networks
Graphs are one of the most versatile data structures. Indeed, vertices and edges can support signals, resulting in a richer structure than tabular data. Common examples are social networks, traffic maps, collaboration networks, and molecules. However, graphs do have their limitations: a graph only permits dyadic interaction between vertices via edges. Multiple recent works have tackle this issue by introducing more general structures, such as simplicial complexes, cellular complexes and hypergraphs [25, 20, 17, 18, 15, 14, 22, 21, 26].
@@ C.1 Simplicial complexes
Simplicial complexes are an extension of graphs, and generalise the notion of triangles by using so-called simplices (simplex in the singular). In general setting, vertices are referred to as 00-simplices, and edges as $1$-simplices. A $2$-simplex is a triangle formed by three adjacent vertices. We observe that this describes a triadic interaction between vertices (a relation between three vertices). We can define simplices of arbitrary dimension $k\geq 0$. However, 00-, $1$- and $2$-simplices are sufficient in numerous interesting situations. A simplicial complex $\mathcal{S}=(\mathcal{V},\mathcal{E},\mathcal{T})$ is constructed from a vertex set $\mathcal{V}$, an edge set $\mathcal{E}$, and a triangle set $\mathcal{T}$. We say that $\mathcal{S}$ is of dimension $2$ and call it a simplicial $2$-complex. Finally, the number of vertices, edges and triangles are denoted by $N_{v},N_{e}$ and $N_{t}$, respectively.
@@
So far, our simplicial complexes are undirected. We can give an orientation to a simplicial complex. An oriented simplicial $2$-complex has an orientation on its edges and triangles, requiring an ordering to be imposed on its vertices. For an edge $e=\{v_{1},v_{2}\}$, an orientation means assigning a direction, transforming $e$ into $(v_{1},v_{2})$ or $(v_{2},v_{1})$. We notice that an edge has always two possible orientations. Giving an orientation amounts to converting an undirected graph into a directed one. Likewise, a triangle $t=\{v_{1},v_{2},v_{3}\}$ has six possible orientations. However, two equivalence classes can be determined, reducing the orientation to be either clockwise or anti-clockwise. We point out that although an orientation is necessary to express concepts such as edge flows, the choice of orientation is arbitrary.
@@
A more general structure is that of cellular complexes. The treatment of cellular complexes is similar to simplicial complexes. For more details, see Alain et al. [15]. For an illustration, we refer to Figure C.1. We could also easily handle hypergraphs by using a Laplacian on hypergraphs.
@@ C.2 Hodge Laplacians
If one takes a step back, it becomes apparent that graphs and simplicial complexes have a chain-like structure: vertices are connected to edges, and edges are connected to triangles. More precisely, we say that vertices have edges as their upper neighbours, edges have vertices as their lower neighbours and triangles as their upper neighbours, and triangles have edges as their lower neighbours. This leads us to the definition of two incidence matrices, $\boldsymbol{B}_{ve}\in\mathbb{Z}^{N_{v}\times N_{e}}$ and $\boldsymbol{B}_{2}\in\mathbb{Z}^{N_{e}\times N_{t}}$. While $\boldsymbol{B}_{ve}$ is the incidence matrix between vertices and edges, $\boldsymbol{B}_{et}$ is the incidence matrix between edges and triangles. In a general setting, these matrices are called boundary matrices, and are denoted by $\boldsymbol{B}_{1}$ and $\boldsymbol{B}_{2}$, respectively, since vertices are $1$-simplices, and edges are $2$-simplices. This results in the general definition of a boundary matrix $\boldsymbol{B}_{k}\in\mathbb{Z}^{N_{k}\times N_{k+1}}$, where $N_{k}$ is the number of $k$-simplices. Note that for a simplicial $K$-complex, $\boldsymbol{B}_{0}=\boldsymbol{0}$ and $\boldsymbol{B}_{k+1}=\boldsymbol{0}$. For simplicial $2$-complexes, it simply means that vertices have no lower neighbours, and triangles have no upper neighbours. We can define a more general notion of Laplacian, called the Hodge Laplacian,
@@
The component $\mathbf{B}_{k}^{\top}\mathbf{B}_{k}$ is sometimes called the lower Laplacian, and $\mathbf{B}_{k+1}\mathbf{B}_{k+1}^{\top}$, the upper Laplacian. We observe that the graph Laplacian is nothing other than $\boldsymbol{L}_{0}$, and the graph Helmholtzian is $\boldsymbol{L}_{1}$.
@@
A $k$-clique is a subset of $k$ vertices, such that every two (distinct) vertices in the clique are adjacent. In graphs, in other words, simplicial $1$-complexes, edges have no upper neighbours. Hence, the edge-to-triangle incidence matrix $\boldsymbol{B}_{2}=\boldsymbol{0}$. In turn, this implies that the graph Helmholtzian is equal to its lower Laplacian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}$. However, it is possible to consider the $3$-cliques of a graph in order to have a non-zero $\boldsymbol{B}_{2}$. Intuitively, this is like considering triangles. We can then have a “full” graph Helmholtzian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}+\boldsymbol{B}_{2}\boldsymbol{B}_{2}^{\top}$.

### v2

@@ 4 Conclusion and Future Directions
We have presented a GP-based classification algorithm for classifying graphs according to one or both vertex and edge features. By applying the graph wavelet transform to spatial graph features, we have constructed spectral features, providing multi-resolution spectral signatures of the original features, subsequently utilised as input points to a standard GP. Furthermore, by taking the discrete Hodge decomposition, we have shown improvements over the method proposed by Opolka et al. [5], even on graph datasets containing only vertex features, owing to our flexible Hodgelet kernel. Overall, we have demonstrated that our approach effectively improves graph classification tasks by employing a spectral perspective to capture both local and global properties of vertex and edge features. In the future, we intend to explore extensions to higher-order networks, including simplicial complexes, cellular complexes, and hypergraphs, which we briefly outline in Appendix C.
@@ Appendix A Gaussian Processes
Gaussian processes are stochastic processes that can be considered as Gaussian distributions extended to infinite dimensions, providing a powerful non-parametric technique for modelling distributions over functions.
More precisely, a Gaussian process over $\mathbb{R}^{d}$ is a random function $f:\Omega\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, characterised by a mean function $\mu:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and a kernel function $\kappa:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, satisfying $\mu(\boldsymbol{x})\coloneq\mathbb{E}[f(\boldsymbol{x})]$ and $\kappa(\boldsymbol{x},\boldsymbol{x}^{\prime})\coloneq\mathrm{Cov}[f(\boldsymbol{x}),f(\boldsymbol{x}^{\prime})]$, respectively, for any $\boldsymbol{x},\boldsymbol{x}^{\prime}\in\mathbb{R}^{d}$. We denote this relation by writing that $f\sim\mathrm{GP}(\mu,\kappa)$. For any finite collection of input points $\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{M}\in\mathbb{R}^{d}$, the random vector $\boldsymbol{f}=[f(\boldsymbol{x}_{1}),\ldots,f(\boldsymbol{x}_{M})]^{\top}\sim\mathrm{N}(\boldsymbol{\mu},\boldsymbol{K})$ is jointly Gaussian, where $\boldsymbol{\mu}=[\mu(x_{1}),\ldots,\mu(x_{M})]^{\top}$ and $(\boldsymbol{K})_{ij}=k(\boldsymbol{x}_{i},\boldsymbol{x}_{j})$.
Gaussian processes are non-parametric in the sense that they define a distribution directly on functions rather than on parameters of a function. This offers the advantage that a Gaussian process can generate functions of increasing sophistication as more data are acquired, unconstrained by a predetermined parametric structure.
@@
A.1 Bayesian inference paradigm
Gaussian processes are employed in the Bayesian inference paradigm:
  1. 1.
    A Gaussian process is defined as a prior distribution over an unknown function, capturing our prior beliefs about the possible shapes and behaviour of this function before observing any data.
  2. 2.
    A likelihood is a modelling choice representing how the functions from the prior distribution generate the data. An essential role of the likelihood is to define the assumptions about the data noise. By observing a dataset, the likelihood assesses the goodness of fit between our prior beliefs and the actual observed data.
  3. 3.
    By combining the prior distribution and the likelihood using the celebrated Bayes’ theorem, a posterior distribution is compute to update our beliefs about the unknown function. For regression, a common assumption is a Gaussian likelihood, leading the posterior distribution to be a Gaussian process again. However, for classification, the likelihood is non-Gaussian, often Bernoulli for binary classification or categorical for multi-class classification. This results in a posterior distribution that is intractable, meaning that it cannot be expressed in a closed-form. In this case, it is possible to derive an approximate posterior distribution using techniques such as variational inference.
  4. 4.
    By conditioning on the dataset, the posterior distribution is employed to derive the predictive posterior distribution for new input points. The mean of this distribution serves as the predictions, while the variance provides a measure of the uncertainty around them.
A.2 Kernel functions
Gaussian process is uniquely defined by a mean function and kernel function, which has the advantage of being simple and interpretable. The mean function is often set to zero, i.e. $\mu(\boldsymbol{x})=0$, for every $\boldsymbol{x}\in\mathbb{R}^{d}$, especially after normalising the dataset since it simplifies derivations and proofs. The kernel function is thus considered the most defining component in describing the behaviour of a Gaussian process and measures the similarities between pairs of input points. It is important to note that a kernel function must be a symmetric and positive semi-definite function as it defines the covariance between two input points. Some popular choices are the square exponential kernel function (also called radial basis function kernel)
@@ A.2 Kernel functions
The parameters of a kernel function are called hyperparameters and they are typically optimised by maximum likelihood estimation.
A.3 Computational cost and memory requirements
@@
Computing the posterior requires inverting a $M\times M$ Gram matrix, where $M$ is the number of training input points. This results to a cubic computational complexity $\mathcal{O}(M^{3})$ and quadratic memory requirements $\mathcal{O}(M^{2})$. Fortunately, sparse Gaussian processes [27] have been devised to alleviate this by constructing a smaller training dataset consisting of $N$ pseudo input points, called inducing points, where $N\ll M$. This technique leads to a computational complexity $\mathcal{O}(MN^{2}+N^{3})$ and memory requirements $\mathcal{O}(MN)$.
Appendix B Spectral Analysis
Spectral analysis is a set of tools for analysing and manipulating signals according their constituent frequencies. These methods have found numerous applications in disciplines as diverse as telecommunications, physics, chemistry, signal processing, quantitative finance and, of course, machine learning.
Suppose a continuous function $f:\mathbb{R}\rightarrow\mathbb{R}$ representing a signal function. This signal function can be expressed as a superposition of sinusoidal plane waves $e^{i2\pi\xi x}$ at different frequencies $\xi$,
[…]
where $\hat{f}(\xi)$ is the Fourier transform of $f$,
[…]
The absolute value $|\hat{f}(\xi)|$ represents the amplitude of $\xi$, reflecting the strength in the original signal of the sinusoidal plane wave associated to $\xi$. The transformed function $\hat{f}$ makes it possible to explore the importance of specific frequencies, revealing insights into patterns, periodicities and trends that may not be apparent from examining the original signal $f$. We observe that this integral transform does not lose information. Indeed, the function $f$ can be recovered from (11) using the inverse Fourier transform (10). The collection of sinusoidal plane waves $\{e^{i2\pi\xi x}\}_{\xi}$ are orthonormal and commonly called the Fourier basis.
Fourier transform and Laplacian.
There are important connections between the Fourier transform and the Laplacian. The Laplacian $\Delta f$ is the divergence of the gradient of $f$,
$\Delta f\coloneq\nabla\cdot\nabla f.$ (12)
It turns out that the Fourier basis function $e^{i2\pi\xi x}$ is the generalised eigenfunction of $\Delta f$ associated to the eigenvalue $\lambda=-(2\pi)^{2}|\xi|^{2}$. We note that small eigenvalues correspond to small frequencies, and large eigenvalues to large frequencies. Consequently, the eigenvalues of the Laplacian are good surrogates for the frequencies.
B.1 Graph wavelet transform
The notion of Fourier transform can be adapted to graphs. Let $\boldsymbol{x}\in\mathbb{R}^{N}$ be a column vector representing a discrete signal on the $N$ vertices of a graph. The eigendecomposition of the graph Laplacian is given by
$\boldsymbol{L}=\boldsymbol{U}\boldsymbol{\Lambda}\boldsymbol{U}^{\top}.$ (13)
Similar to the continuous case, the eigenvector matrix $\boldsymbol{U}$ is referred to as the graph Fourier basis, and we consider the eigenvalues of $\boldsymbol{L}$ as the graph frequencies.
The graph Fourier transform $\boldsymbol{U}^{T}$ performs a projection of the signal onto the graph Fourier basis, thus producing a finite number of Fourier coefficients
$\hat{\boldsymbol{x}}\coloneq\boldsymbol{U}^{\top}\boldsymbol{x}\in\mathbb{R}^{N}.$ (14)
We see that the Fourier coefficients belong to the eigenspace of the graph Laplacian. They provide complete localisation in terms of frequency, meaning that every frequency is represented and its contribution to the original signal can be identified. By contrast, the original signal offers complete resolution in space (or time), making it possible to determine how the signal varies across the vertices. However, it is often advantageous not to be limited to just one or the other. The spectral graph wavelet transform offers multi-scale resolution, allowing a more balanced analysis between the spatial and frequency domains. Instead of sinusoidal plane waves, it employs the more general notion of wavelets. A wavelet can be scaled and translated, meaning that it can be tuned to capture localised changes in space. By contrast, a sinusoidal plane wave is uniform and extend across the entire signal.
The idea behind wavelet transform is to derive the wavelet coefficients from the Fourier coefficients by applying a wavelet filter $w:\mathbb{R}\rightarrow\mathbb{R}$ on the eigenvalues of the graph Laplacian and then performing the inverse Fourier transform,
$\hat{\boldsymbol{x}}\coloneq\boldsymbol{U}w(\boldsymbol{\Lambda})\boldsymbol{U}^{\top}\boldsymbol{x}.$ (15)
A wavelet filter is a combination of a scaling function $a:\mathbb{R}\rightarrow\mathbb{R}$ at the scale $\alpha$ and a wavelet function $b:\mathbb{R}\rightarrow\mathbb{R}$ at $L$ different scales $\beta_{l}$,
$w(\lambda)\coloneq a(\alpha\lambda)+\sum_{l=1}^{L}b(\beta_{l}\lambda),\quad\quad\Theta\coloneq\{\alpha,\beta_{1},\dots,\beta_{L}\}.$ (16)
The wavelet function at each scale represents a scaled and translated variant of a mother wavelet. The role of the wavelet function is to serve as a band-pass filter, covering medium and high frequencies. The scaling function is a low-pass filter, capturing low frequencies.
@@ B.1 Graph wavelet transform
The reader is referred to Hammond et al. [8] for a more detailed description of the graph wavelet transform.
Mother wavelet.
@@ Appendix C Extension to Higher-Order Networks
Graphs can support signals on their vertices and edges, resulting in a richer structure than tabular data. Common examples are social networks, electrical grids, citation networks, traffic maps, chemical reaction networks, collaboration networks, and molecules. However, graphs do have an important limitation: a graph only permits dyadic interactions between its vertices via its edges. Multiple recent works have tackle this issue by focusing on more general structures, such as simplicial complexes, cellular complexes and hypergraphs [28, 22, 18, 19, 16, 15, 24, 23, 29]. For a comparative illustration between a graph, simplicial complex and a cellular complex, we refer to Figure C.1.
@@ C.1 Simplicial complexes
Simplicial complexes represent a natural extension of graphs, and generalise the notion of vertices and edges by using simplices or simplex in the singular. In this broader setting, vertices are referred to as 00-simplices, and edges as $1$-simplices. A $2$-simplex is a triangle formed by three adjacent vertices. We observe that this describes a triadic interaction between vertices, in other words, a relation between three vertices. We can create simplices of arbitrary dimension $k\geq 0$. However, the first three dimensions are often sufficient. We define a simplicial complex $\mathcal{S}=(\mathcal{V},\mathcal{E},\mathcal{T})$ has a structure constructed from a vertex set $\mathcal{V}$, an edge set $\mathcal{E}$, and a triangle set $\mathcal{T}$. We say that $\mathcal{S}$ is a simplicial $2$-complex or a simplicial complex of dimension $2$. Finally, the number of vertices, edges, and triangles are denoted by $N_{v},N_{e}$ and $N_{t}$, respectively.
@@
So far, our simplicial complexes are undirected. However, we can give an orientation to a simplicial complex. An oriented simplicial $2$-complex has an orientation on its edges and triangles, requiring an ordering to be imposed on its vertices. For an edge $e=\{v_{1},v_{2}\}$, an orientation means assigning a direction, transforming $e$ into either $(v_{1},v_{2})$ or $(v_{2},v_{1})$. We notice that an edge has always exactly two possible orientations. Likewise, a triangle $t=\{v_{1},v_{2},v_{3}\}$ has six possible orientations. However, two equivalence classes can be determined, reducing the orientation to be either clockwise or anti-clockwise. We point out that although an orientation is necessary to express concepts such as edge flows, the choice of orientation is arbitrary. Giving an orientation seems like converting an undirected graph into a directed one, but it is fundamentally different.
@@
A more general structure is that of cellular complexes. The treatment of cellular complexes is similar to simplicial complexes. For more details, see Alain et al. [16]. We could also easily handle hypergraphs by using a Laplacian on hypergraphs.
@@ C.2 Hodge Laplacians
If one takes a step back, it becomes apparent that graphs and simplicial complexes have a chain-like structure: vertices are connected to edges and edges are connected to triangles. More precisely, we say that vertices have edges as their upper neighbours, edges have vertices as their lower neighbours and triangles as their upper neighbours, and triangles have edges as their lower neighbours. This leads us to the definition of two incidence matrices, $\boldsymbol{B}_{ve}\in\mathbb{Z}^{N_{v}\times N_{e}}$ and $\boldsymbol{B}_{et}\in\mathbb{Z}^{N_{e}\times N_{t}}$, sometimes called vertex-to-edge and edge-to-triangle incidence matrices, respectively. While $\boldsymbol{B}_{ve}$ is the incidence matrix between vertices and edges, $\boldsymbol{B}_{et}$ is the incidence matrix between edges and triangles. In a more general setting, these matrices are simply called boundary matrices and they are denoted by $\boldsymbol{B}_{1}$ and $\boldsymbol{B}_{2}$, respectively, since vertices are $1$-simplices and edges are $2$-simplices. This results in the general definition of a boundary matrix $\boldsymbol{B}_{k}\in\mathbb{Z}^{N_{k}\times N_{k+1}}$, where $N_{k}$ is the number of $k$-simplices. Note that for a simplicial $K$-complex, $\boldsymbol{B}_{0}=\boldsymbol{0}$ and $\boldsymbol{B}_{k+1}=\boldsymbol{0}$. This simply means that 00-simplices have no lower neighbours, and $k$-simplices have no upper neighbours. We can define a more general notion of Laplacian, called the Hodge Laplacian,
@@
The component $\mathbf{B}_{k}^{\top}\mathbf{B}_{k}$ is sometimes called the lower Laplacian, and $\mathbf{B}_{k+1}\mathbf{B}_{k+1}^{\top}$, the upper Laplacian. We observe in particular that the graph Laplacian is nothing other than $\boldsymbol{L}_{0}$, and the graph Helmholtzian is $\boldsymbol{L}_{1}$.
Interested readers are invited to find out more by looking at the concept of chain complex.
@@
A $k$-clique is a subset of $k$ vertices, such that every two distinct vertices in the clique are adjacent. Intuitively, this is like considering triangles. Edges have no upper neighbours in simplicial $1$-complexes, i.e. graphs. Hence, the edge-to-triangle incidence matrix $\boldsymbol{B}_{2}=\boldsymbol{0}$. In turn, this implies that the graph Helmholtzian is equal to its lower Laplacian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}$. Nevertheless, we can consider the triangles of a graph in order to have a non-zero $\boldsymbol{B}_{2}$. We can then have a complete graph Helmholtzian, i.e. $\boldsymbol{L}_{1}=\boldsymbol{B}_{1}^{\top}\boldsymbol{B}_{1}+\boldsymbol{B}_{2}\boldsymbol{B}_{2}^{\top}$.